OpenStudy (anonymous):

Can't get the right answer for some reason! The second angle of a triangle is four less than the first, and the third angle is six times the first. Find the measures of the angles if the angles in a triangle all add up to 180 degrees.

OpenStudy (anonymous):

I already know to set up the problem like this x+x-4+6x =180

OpenStudy (oaktree):

That's right so far. So what's your next step?

OpenStudy (oaktree):

Well, first off, do you get 23 for the first, 19 for the second, and 138 for the third?

OpenStudy (oaktree):

@DimDanny?

OpenStudy (anonymous):

Yes, I know the answers but, i keep getting 26, 22, and 156. Can you walk me through how you get the correct answer?

OpenStudy (oaktree):

Okay - now I got it. So what do you do after you have the equation? I don't want to just give you how to do it - it helps if you figure it out yourself.

OpenStudy (anonymous):

ok, well i cancel out the two x's and get -4+6x=180 is that correct or do i combine?

OpenStudy (oaktree):

Combine. You wouldn't cancel unless one was positive and one was negative, and they're both positive. You should get 8x-4=180.

OpenStudy (anonymous):

OH! so now, i add 4 to both sides then divide by 8 both sides correct?

OpenStudy (oaktree):

Yup.

OpenStudy (anonymous):

Thanks I'll probably post more of these questions because im trying to study for a test.

OpenStudy (oaktree):

Okay. I'm switching between you and someone else right now, so forgive me if I'm slow.